OpenStudy (anonymous):

Help please will give a medal. A pizza baker uses 2.5 cups of flour to make each large pizza. If f represents the number of cups of flour, which equation models this situation? A) f= 2.5p, where p is 1 large pizza B) f= 2.5p, where p is the number of large pizzas C) f= p/2.5, where p is the number of cups of flour in a large pizza D) f= p/2.5, where p is the number of large pizzas

OpenStudy (anonymous):

@TBNRfrags

OpenStudy (anonymous):

it will be a

OpenStudy (anonymous):

i know it will be A) f= 2.5p, where p is 1 large pizza

OpenStudy (anonymous):

A was not correct.

OpenStudy (anonymous):

its B

OpenStudy (gorv):

total number of cup=f*2.5

OpenStudy (gorv):

p=f*2.5

OpenStudy (gorv):

f=p/2.5

OpenStudy (tbnrfrags):

gorv is correct ^

OpenStudy (gorv):

D

OpenStudy (anonymous):

Thank you @gorv and @TBNRfrags

OpenStudy (tbnrfrags):

np :)